OpenStudy (anonymous):

Give the best method for solving 2y^2+8y=0 a.factoring b.property of sqaure roots of equal numbers c.completing the square d. quadratic formula

OpenStudy (amistre64):

there is no "best method"

OpenStudy (anonymous):

welll its in my book and its multiple choice and those are the answers so i have to pick one

OpenStudy (amistre64):

the "best method" is the one that gets a solution. and since b is the only thing i dont recognize id try to avoid that one

OpenStudy (anonymous):

so quadratic formula

OpenStudy (anonymous):

factoring is the easiest..

OpenStudy (anonymous):

I'd say factoring...

OpenStudy (amistre64):

by factoring: 2y^2 + 8y = 0 2y (y + 4) = 0 y = 0, -4 .......................... by CTS 2y^2 + 8y = 0 y^2 + 4y = 0 y^2 + 4y +4 = 4 (y+2)^2 = 4 y+2 = +- sqrt(4) y = -2 +- 2 ; y=0, -4 ------------------------ quad form 2y^2 + 8y = 0 -8 +- sqrt(8^2-4(2)(0)) -------------------- 2(2) -8 +- 8 ------- ; y = 0, -4 4

OpenStudy (amistre64):

there is no "best" method of solving this

OpenStudy (anonymous):

k thanks youn tried but you stilll arent getting what im asking but whatever